# Second- and third-order properties of multidimensional Langevin equations

### v12

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
upon dividing by $\tau$ and taking $\tau\to 0$, where we used $\langle\mathbf{x}(\tau)\mathbf{x}(\tau)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Eq. (5) follows if the derivative can be taken inside the expectation (however, see Appendix B for a counterexample). This type of manipulation will become useful later.
@@ 2.2. Comparison of experimental data with theoretical predictions
Now if we neglect permutations in $(i,j)$ and $(i^{\prime},j^{\prime})$ [^5] I.e., only one of $(i,j)$ and $(j,i)$ is included in the list of “super-indices” (and similarly for $(i^{\prime},j^{\prime})$). This is necessary to invert $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$, since this quantity is invariant upon swapping $i\leftrightarrow j$ (or $i^{\prime}\leftrightarrow j^{\prime}$)., we can see that the matrix product of $M^{ij}M^{i^{\prime}j^{\prime}}$ with the matrix inverse of $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$ has left eigenvector $v_{ij}$ and eigenvalue $\mu$. Thus we may consider instead of Eq. (10) the relation:
@@ 2.3. Probability current density and angular momentum
(30) $\mathbf{L}:=\lim_{\tau\to 0}\frac{\langle\mathbf{x}(0)\mathbf{x}(\tau)^{\mathsf{T}}-\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle}{\tau}=\langle\mathbf{x}\dot{\mathbf{x}}^{\mathsf{T}}-\dot{\mathbf{x}}\mathbf{x}^{\mathsf{T}}\rangle=\mathbf{C}\mathbf{A}^{\mathsf{T}}-\mathbf{A}\mathbf{C},$
@@
Now we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. The angular momentum is $L(x,y):=\langle x\dot{y}-\dot{x}y\rangle$ (where the time-derivatives are interpreted in Itô sense). The angular momentum is related to the matrix $\bm{\Omega}$ by the relation
(31) $\mathbf{A}\mathbf{C}-\mathbf{C}\mathbf{A}^{\mathsf{T}}=\begin{pmatrix}0&-L\\ L&0\end{pmatrix}=2\bm{\Omega}\mathbf{C},$
and upon taking determinants, we get
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
(64) $\mathbf{L}(\mathbf{x},y):=\lim_{\tau\to 0}\frac{\langle(\mathbf{x}(0)+\mathbf{x}(\tau))(y(\tau)-y(0))\rangle}{\tau}$
@@
The condition for detailed balance is $\mathbf{L}(\mathbf{x},y)=\mathbf{0}$ and $L(x^{i},x^{j})=0$ for all $i,j$.
@@ 3.3. Comparison of experimental data with theoretical predictions
(84) $2D^{yz}=\lim_{\tau\to 0}\frac{\langle(y(\tau)-y(0))(z(\tau)-z(0))\rangle}{\tau}=\text{``$\langle-\dot{y}z-y\dot{z}\rangle$''}.$ ” .
Since there is no angular momentum between $y$ and $z$, we assign $\dot{y}z$ and $y\dot{z}$ the same (fictional) value, $-D^{yz}$. For non-zero time lags, we integrate:
@@ 4. Third-order properties
where $\mathbf{A}$, $\mathbf{D}$, $a^{i}_{jk}$, and $b^{ij}_{k}$ are constant coefficients, we have suppressed $t$-dependence on the r.h.s., and Einstein summation notation is used. Neither addition to the model can stand on its own: in general, $a^{i}_{jk}$ by itself leads to divergences of $\mathbf{x}\to\infty$, and $b^{ij}_{k}$ by itself renders $\langle\mathrm{d}[x^{i},x^{j}]/\mathrm{d}t\mid\mathbf{x}(t)\rangle$ not positive semidefinite. A third possible contribution,
(92) $\lim_{\tau\to 0}\frac{\langle(x^{i}(\tau)-x^{i}(0))(x^{j}(\tau)-x^{j}(0))(x^{k}(\tau)-x^{k}(0))\rangle}{\tau}=0,$
@@
(98) $\widetilde{L}(x^{i},x^{j},x^{k}):=\lim_{\tau\to 0}\frac{\langle x^{i}(\tau/2)(x^{j}(0)x^{k}(\tau)-x^{j}(\tau)x^{k}(0))\rangle}{\tau}.$
@@
(103) $\lim_{\tau\to 0}\frac{\langle(x^{i}(0)x^{j}(\tau)+x^{i}(\tau)x^{j}(0))(y(\tau)-y(0))\rangle}{\tau}=\lim_{\tau\to 0}\frac{\langle x^{i}(\tau/2)(x^{j}(0)+x^{j}(\tau))(y(\tau)-y(0))\rangle}{\tau}=L(x^{i}x^{j},y).$
@@
(104) $\lim_{\tau\to 0}\frac{\langle x^{i}(\tau)(x^{j}(\tau)-x^{j}(0))(x^{k}(\tau)-x^{k}(0))\rangle}{\tau}=\lim_{\tau\to 0}\frac{\langle x^{i}(0)(x^{j}(\tau)-x^{j}(0))(x^{k}(\tau)-x^{k}(0))\rangle}{\tau}.$
@@ 5. Inhomogeneous diffusion
(105) $0=\langle(\mathbf{A}\mathbf{x})^{i}x^{j}x^{k}\rangle+\langle x^{i}(\mathbf{A}\mathbf{x})^{j}x^{k}\rangle+\langle x^{i}x^{j}(\mathbf{A}\mathbf{x})^{k}\rangle+2b^{ij}_{l}C^{kl}+2b^{ik}_{l}C^{jl}+2b^{jk}_{l}C^{il},$
@@
(123) $\left\langle\frac{\mathrm{d}[y,y](t)}{\mathrm{d}t}\mid\mathbf{x}(t)\right\rangle$ $=2D^{yy}+2b^{yy}_{i}x^{i}.$
Then:
@@ 6.5. Integrated variables
We now address the case of an integrated variable $y$, where
(174) $\langle\dot{y}\mid\mathbf{x}\rangle=A^{y}_{i}x^{i}+a^{y}_{ij}(x^{i}x^{j}-C^{ij}).$
We can define:
@@ 6.8. Quantitative significance of third-order effects
Following the discussion of deviations of experiment from theory for the second moments, we may treat $(i,j,k)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime},k^{\prime})$, then multiply the l.h.s. by the matrix inverse of the r.h.s. For the same reason as before, the resulting matrix has a single non-zero eigenvalue $\mu$ equal to its trace. To satisfy symmetry, we may consider an ensemble of stochastic systems whose “ensemble covariance” of third moments (l.h.s. of the above) is:
@@ 7.3. Inhomogeneous diffusion
We also need for $n\geq 1$:
@@
The base case $n=1$ for Eq. (248) is an assumption needed for time-reversal symmetry. The base case $n=1$ for Eq. (249) is a consequence of Eq. (246) and Eq. (247), as the latter implies:
@@
Again, the base case $n=1$ of Eq. (254) is a consequence of Eqs. (246) and (252), while Eq. (253)–(254) can be shown by mathematical induction on $n$. This establishes the time-reversibility of $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$.
@@
Once again, the case $n=1$ for Eq. (257) is a consequence of Eqs. (246) and (252), while Eq. (256)–(257) can be shown assuming the case $n=1$. This establishes the time-reversibility of $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$.
@@ 7.4. Nonlinear drift
Again, the base case $n=1$ for Eq. (272) follows from Eqs. (246)–(247). To carry out the proof by induction for Eq. (271)–(272) assuming the base case $n=1$, we need to use, for $i\geq 1$ and $n\geq 1$:
@@ 7.5. Integrated variables
where the base case $n=1$ for Eq. (284) follows from $L(\mathbf{x}^{(2)+},y)=0$ and for Eq. (285) follows from Eqs. (246) and (252). This shows the time-reversibility of the covariance function $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. For $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$, we need to show that for $n\geq 1$:
@@
where the base case $n=1$ for the upper sign follows from $L(\mathbf{x}^{(2)+},y)=0$ and for the lower sign follows from Eqs. (246) and (252). Here, we need to use:
@@ 7.6. Second-order covariance functions to quadratic order
where the base case $n=1$ is assumed for Eq. (292), and for Eq. (293) follows from our assumptions that $\mathbf{D}^{+-}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$ are $a^{2}\mathcal{O}((a,b)^{2})$, and that $\langle\mathbf{x}^{(2)\pm}{\mathbf{x}^{\mp}}^{\mathsf{T}}\rangle=a\mathcal{O}((a,b)^{2})$. Using the properties stated in section 7.4, the induction step is readily performed. For the case of an integrated variable $y$ of even parity, we additionally need the equalities in section 7.5 and:
@@ 8.2. Almost Markovian dynamics
However, the expression on the r.h.s. results in a non-differentiable covariance function at $\tau=0$. The differentiability at $\tau=0$ is expressed as:
@@ 8.3. Third-order properties
(327) $\left\langle\frac{\mathrm{d}[v^{i},v^{j}](t)}{\mathrm{d}t}\mid\mathbf{x}(t),\mathbf{v}(t)\right\rangle$ $=2\mu D^{ij}+2\mu(b_{x})^{ij}_{k}x^{k}+2\mu(b_{v})^{ij}_{k}v^{k}.$
We first note that the estimation of quantities $\langle f(\mathbf{x},\mathbf{v})\dot{\mathbf{v}}\rangle$ from measured time-series data $\mathbf{x}(n\Delta t)$, $n=0,1,2,\ldots$ is not what might be naïvely expected from time-discretization of $\mathbf{v}$ and $\dot{\mathbf{v}}$, but has a correction due to fluctuations [10]. Aside from that, there is the issue of measuring quantities that are robust against measurement noise, for which the authors of Ref. [10] consider quantities of the form:
@@
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (93)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\mathcal{B}$ and measurement error covariance $\bm{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) in [10] should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in their code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\bm{\eta}$, where $\bm{\eta}$ is measurement noise. Using the second differences:

### v13

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
upon dividing by $\tau>0$ and taking $\tau\to 0^{+}$, where we used $\langle\mathbf{x}(\tau)\mathbf{x}(\tau)^{\mathsf{T}}\rangle=\langle\mathbf{x}(0)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Eq. (5) follows if the derivative can be taken inside the expectation (however, see Appendix B for a counterexample). This type of manipulation will become useful later.
@@ 2.2. Comparison of experimental data with theoretical predictions
Now if we neglect permutations in $(i,j)$ and $(i^{\prime},j^{\prime})$ [^5] I.e., only one of $(i,j)$ and $(j,i)$ is included in the list of “super-indices” (and similarly for $(i^{\prime},j^{\prime})$). This is necessary to invert $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$, since this quantity is invariant upon swapping $i\leftrightarrow j$ (or $i^{\prime}\leftrightarrow j^{\prime}$)., we can see that the matrix product of $M^{ij}M^{i^{\prime}j^{\prime}}$ with the matrix inverse of $(C^{ii^{\prime}}C^{jj^{\prime}}+C^{ij^{\prime}}C^{ji^{\prime}})/2$ has left eigenvector with elements $v_{ij}+v_{ji}$ for $i\neq j$ and $v_{ij}$ for $i=j$ and eigenvalue $\mu$. Thus we may consider instead of Eq. (10) the relation:
@@ 2.3. Probability current density and angular momentum
(30) $\mathbf{L}:=\lim_{\tau\to 0^{+}}\frac{\langle\mathbf{x}(0)\mathbf{x}(\tau)^{\mathsf{T}}-\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle}{\tau}=\langle\mathbf{x}\dot{\mathbf{x}}^{\mathsf{T}}-\dot{\mathbf{x}}\mathbf{x}^{\mathsf{T}}\rangle=\mathbf{C}\mathbf{A}^{\mathsf{T}}-\mathbf{A}\mathbf{C},$
@@
Now we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. The angular momentum is $L:=\langle x\dot{y}-\dot{x}y\rangle$. It is related to the matrix $\bm{\Omega}$ by the relation
(31) $\begin{pmatrix}0&-L\\ L&0\end{pmatrix}=\mathbf{A}\mathbf{C}-\mathbf{C}\mathbf{A}^{\mathsf{T}}=2\bm{\Omega}\mathbf{C},$
where we used the Lyapunov relation. Upon taking determinants, we get
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
(64) $\mathbf{L}(\mathbf{x},y):=\lim_{\tau\to 0^{+}}\frac{\langle(\mathbf{x}(0)+\mathbf{x}(\tau))(y(\tau)-y(0))\rangle}{\tau}$
@@
The condition for detailed balance is $\mathbf{L}(\mathbf{x},y)=\mathbf{0}$ and $L(x^{i},x^{j})=0$ for all $i,j$, where we used the notation $L(x^{i},x^{j}):=\langle x^{i}\dot{x}^{j}-\dot{x}^{i}x^{j}\rangle$.
@@ 3.3. Comparison of experimental data with theoretical predictions
(84) $2D^{yz}=\lim_{\tau\to 0^{+}}\frac{\langle(y(\tau)-y(0))(z(\tau)-z(0))\rangle}{\tau}=\text{``$\langle-\dot{y}z-y\dot{z}\rangle$''}.$ ” .
Since there is no angular momentum between $y$ and $z$, we assign $\langle\dot{y}z\rangle$ and $\langle y\dot{z}\rangle$ the same (fictional) value, $-D^{yz}$. For non-zero time lags, we integrate:
@@ 4. Third-order properties
where all coefficients are constant, we have suppressed $t$-dependence on the r.h.s., and Einstein summation notation is used. Neither addition to the model can stand on its own: in general, $a^{i}_{jk}$ by itself leads to divergences of $\mathbf{x}\to\infty$, and $b^{ij}_{k}$ by itself renders $\langle\mathrm{d}[x^{i},x^{j}]/\mathrm{d}t\mid\mathbf{x}(t)\rangle$ not positive semidefinite. A third possible contribution,
(92) $\lim_{\tau\to 0^{+}}\frac{\langle(x^{i}(\tau)-x^{i}(0))(x^{j}(\tau)-x^{j}(0))(x^{k}(\tau)-x^{k}(0))\rangle}{\tau}=0,$
@@
(98) $\widetilde{L}(x^{i},x^{j},x^{k}):=\lim_{\tau\to 0^{+}}\frac{\langle x^{i}(\tau/2)(x^{j}(0)x^{k}(\tau)-x^{j}(\tau)x^{k}(0))\rangle}{\tau}.$
@@
(103) $\lim_{\tau\to 0^{+}}\frac{\langle(x^{i}(0)x^{j}(\tau)+x^{i}(\tau)x^{j}(0))(y(\tau)-y(0))\rangle}{\tau}=\lim_{\tau\to 0^{+}}\frac{\langle x^{i}(\tau/2)(x^{j}(0)+x^{j}(\tau))(y(\tau)-y(0))\rangle}{\tau}=L(x^{i}x^{j},y).$
@@
(104) $\lim_{\tau\to 0^{+}}\frac{\langle x^{i}(\tau)(x^{j}(\tau)-x^{j}(0))(x^{k}(\tau)-x^{k}(0))\rangle}{\tau}=\lim_{\tau\to 0^{+}}\frac{\langle x^{i}(0)(x^{j}(\tau)-x^{j}(0))(x^{k}(\tau)-x^{k}(0))\rangle}{\tau}.$
@@ 5. Inhomogeneous diffusion
(105) $0=\frac{\mathrm{d}}{\mathrm{d}t}\langle x^{i}x^{j}x^{k}\rangle=\langle(\mathbf{A}\mathbf{x})^{i}x^{j}x^{k}\rangle+\langle x^{i}(\mathbf{A}\mathbf{x})^{j}x^{k}\rangle+\langle x^{i}x^{j}(\mathbf{A}\mathbf{x})^{k}\rangle+2b^{ij}_{l}C^{kl}+2b^{ik}_{l}C^{jl}+2b^{jk}_{l}C^{il},$
@@
(123) $\left\langle\frac{\mathrm{d}[y,y](t)}{\mathrm{d}t}\mid\mathbf{x}(t)\right\rangle$ $=2D^{yy}+2b^{yy}_{i}x^{i},$
where all coefficients are constant. Then:
@@ 6.5. Integrated variables
We now address the case of an integrated variable $y$ obeying:
(174) $\langle\dot{y}\mid\mathbf{x}\rangle=A^{y}_{i}x^{i}+a^{y}_{ij}(x^{i}x^{j}-C^{ij}),$
where all coefficients are constant. We can define:
@@ 6.8. Quantitative significance of third-order effects
Following the discussion of deviations of experiment from theory for the second moments, we may treat $(i,j,k)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime},k^{\prime})$, then multiply the l.h.s. by the matrix inverse of the r.h.s. For the same reason as before, the resulting matrix has a single non-zero eigenvalue $\mu$ equal to its trace. To satisfy symmetry, using the same argument as before, we may consider an ensemble of stochastic systems whose “ensemble covariance” of third moments (l.h.s. of the above) is:
@@ 7.3. Inhomogeneous diffusion
where the signs denote the parities of the corresponding variables. We also need for $n\geq 1$:
@@
The base case $n=1$ for Eq. (248) is an assumption needed for time-reversal symmetry. The base case $n=1$ for Eq. (249) follows from differentiating Eq. (246) and using Eq. (247), as the latter implies:
@@
Similarly as before, the base case $n=1$ of Eq. (253) is an assumption needed for time-reversal symmetry, while the base case $n=1$ of Eq. (254) is a consequence of Eq. (252), as:
(255) $\lim_{\tau\to 0^{+}}\frac{\langle\mathbf{x}^{-}(0)\mathbf{x}^{+}(\tau)^{\mathsf{T}}(y(\tau)-y(0))\rangle}{\tau}=\lim_{\tau\to 0^{+}}\frac{\langle\mathbf{x}^{-}(\tau)\mathbf{x}^{+}(0)^{\mathsf{T}}(y(\tau)-y(0))\rangle}{\tau}=\langle\mathbf{x}^{-}{\mathbf{x}^{+}}^{\mathsf{T}}\dot{y}\rangle.$
Eq. (253)–(254) can then be shown by mathematical induction on $n$. This establishes the time-reversibility of $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$.
@@
Again similarly, the case $n=1$ for Eq. (257) is an assumption needed for time-reversal symmetry, while the case $n=1$ for Eq. (258) is a consequence of Eqs. (252) and (246). Eq. (257)–(258) can be shown by mathematical induction on $n$. This establishes the time-reversibility of $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$.
@@ 7.4. Nonlinear drift
Similar as before, the base case $n=1$ for Eq. (272) is an assumption needed for time-reversal symmetry, while the base case $n=1$ for Eq. (273) follows from Eqs. (246)–(247). To carry out the proof by induction for Eq. (272)–(273), we need to use, for $i\geq 1$ and $n\geq 1$:
@@ 7.5. Integrated variables
where the base case $n=1$ for Eq. (285) follows from assuming $L(\mathbf{x}^{(2)+},y)=0$ and using Eq. (103), while the base case $n=1$ for Eq. (286) follows from Eq. (252). This shows the time-reversibility of the covariance function $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. For $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$, we need to show that for $n\geq 1$:
@@
where the base case $n=1$ for the upper sign follows from assuming $L(\mathbf{x}^{(2)+},y)=0$ and for the lower sign follows from Eq. (252) and (246). Here, we need to use:
@@ 7.6. Second-order covariance functions to quadratic order
where the base case $n=1$ is assumed for Eq. (293), and for Eq. (294) follows from differentiating $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle$. Using the properties stated in section 7.4, the induction step is readily performed. For the case of an integrated variable $y$ of even parity, we additionally need the equalities in section 7.5 and:
@@ 8.2. Almost Markovian dynamics
However, the expression on the r.h.s. results in a covariance function whose derivative at $\tau=0$ is not antisymmetric. The antisymmetry of the derivative of the covariance function at $\tau=0$ is expressed as:
@@ 8.3. Third-order properties
(328) $\left\langle\frac{\mathrm{d}[v^{i},v^{j}](t)}{\mathrm{d}t}\mid\mathbf{x}(t),\mathbf{v}(t)\right\rangle$ $=2\mu D^{ij}+2\mu(b_{x})^{ij}_{k}x^{k}+2\mu(b_{v})^{ij}_{k}v^{k},$
where all coefficients are constant. We first note that the estimation of quantities $\langle f(\mathbf{x},\mathbf{v})\dot{\mathbf{v}}\rangle$ from measured time-series data $\mathbf{x}(n\Delta t)$, $n=0,1,2,\ldots$ is not what might be naïvely expected from time-discretization of $\mathbf{v}$ and $\dot{\mathbf{v}}$, but has a correction due to fluctuations [10]. Aside from that, there is the issue of measuring quantities that are robust against measurement noise, for which the authors of Ref. [10] consider quantities of the form:
@@
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (93)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\bm{\mathcal{B}}$ and measurement error covariance $\bm{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) in [10] should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in their code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\bm{\eta}$, where $\bm{\eta}$ is measurement noise. Using the second differences: